# Optimal low-rank approximations for linear Gaussian inverse problems on Hilbert spaces Part I: posterior covariance approximation

### v3

@@ 1.4 Outline
We introduce key notation in Section 1.5 below. In Section 2, we begin by recalling the infinite-dimensional formulation of the linear Gaussian inverse problem, and formulate the posterior covariance and posterior mean approximation classes that define our measure approximation problems. In Section 3 we recall the Feldman–Hajek theorem, which characterises when two Gaussian measures are equivalent, and recall expressions for the Kullback–Leibler divergence and Rényi divergence of two equivalent Gaussians. We state the first main result of this paper, Proposition 3.7, which identifies the generalised eigenpairs of the three operator pencils mentioned in the introduction and identifies the finite-dimensional subspace on which the posterior differs from the prior. In Section 4, we consider measure approximation problems where the posterior covariance is approximated and identify solutions in Theorem 4.21 and Corollary 4.23. Auxiliary results are presented in Appendix A, and proofs of the results in this work can be found in Appendix B.
@@ 1.5 Notation
For $I$ a non-empty interval in $\mathbb{R}$, $\ell^{2}(I)$ denotes the space of square-summable sequences, i.e. $\ell^{2}(I)=\{(x_{i})_{i\in\mathbb{N}}\subset I\ :\ \sum_{i\in\mathbb{N}}\lvert x_{i}\rvert^{2}<\infty\}$. If $I\subset\mathbb{R}$ open, then $C^{1}(I)$ denotes the continuously differentiable functions on $I$.
@@ 2 Low-rank posterior covariance approximations
see e.g. [32, Example 6.23]. Equation (3c) should be understood to imply the following two facts: $\operatorname{ran}{\mathcal{C}}_{\textup{pos}}\coloneqq\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}+H}=\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$, and $\mathcal{C}_{\textup{pr}}^{-1}+H:\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\rightarrow\mathcal{H}$ is the inverse of (3b). While all nondegenerate Gaussians are equivalent in a finite-dimensional setting, this is no longer true in an infinite-dimensional setting, where in fact it holds that nondegenerate Gaussians that are not equivalent must be mutually singular. By [32, Theorem 6.31], $\mu_{\textup{pos}}$ and $\mu_{\textup{pr}}$ are in fact equivalent. In particular, $\mu_{\textup{pos}}$ is a nondegenerate measure and the above properties of $\mathcal{C}_{\textup{pr}}$ also hold for $\mathcal{C}_{\textup{pos}}$. We shall construct approximations to $\mu_{\textup{pos}}$ that are equivalent to $\mu_{\textup{pos}}$.
@@ 4.2 Equivalence to target measures of low-rank Gaussian approximations
A natural generalisation to infinite dimensions of the low-rank approximation classes for covariance and precision considered in the finite-dimensional setting of [31, eqs. (2.4) and (4.1)], is to take $\widetilde{\mathscr{C}}_{r}\coloneqq\{\mathcal{C}_{\textup{pr}}-KK^{*}>0:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$ and $\widetilde{\mathscr{P}}_{r}=\{\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}:\ U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$. Let $\mathscr{C}_{r,\mathcal{E}}\coloneqq\{\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathcal{E}:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$ as in the start of Section 4.2. By the preceding discussion, we have the proper inclusions $\widetilde{\mathscr{P}}_{r}^{-1}\subset\mathscr{C}_{r,\mathcal{E}}$, and by the equivalence (iii)(b)$\Leftrightarrow$(iii)(a) of Lemma 4.5(iii) also the proper inclusion $\mathscr{C}_{r,\mathcal{E}}\subset\widetilde{\mathscr{C}}_{r}$ holds. That is, in general $\widetilde{\mathscr{P}}_{r}^{-1}$ contains strictly fewer covariances than those that maintain equivalence between the resulting approximate and exact posterior, while $\widetilde{\mathscr{C}}_{r}$ constains strictly more. For the loss classes considered in this work which require densities to exist, the set $\widetilde{\mathscr{C}}_{r}$ is thus not suitable. Note that by the definition of $\mathscr{P}_{r}$ in (5), $\widetilde{\mathscr{P}}_{r}=\mathscr{P}_{r}$. The fact that $\widetilde{\mathscr{P}}_{r}^{-1}\subset\mathscr{C}_{r,\mathcal{E}}$ motivates the use of approximation class $\widetilde{\mathscr{P}}_{r}$ in this work. Because the form of the precision updates in $\widetilde{\mathscr{P}}_{r}$ parallels the form in [31, eq. (4.1)] that lies at the core of the approximation procedure of [31], our work can be considered to naturally generalise [31].
@@ 4.3 Differentiability and minimisers of covariance loss function
In order to solve Problem 4.4, we formulate it as a minimisation problem over the set of $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. By Corollary 4.8(ii), solving Problem 4.3 is equivalent to solving Problem 4.3. We want to find the minimiser of the function
@@
We finally define the function $g:\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\rightarrow\mathcal{B}(\mathcal{H})_{\mathbb{R}}$ by
@@
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. If $\mathcal{W}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$ is a subspace of finite dimension that contains $g(U+tV)$ for all $t\in\mathbb{R}$, then the restriction ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}:\mathcal{W}\rightarrow\mathbb{R}$ of $F_{f}\circ\Lambda$ to $\mathcal{W}$ satisfies ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)=F_{f}\circ\Lambda\circ g(U+tV)$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g$ is Gateaux differentiable at $U$ in the direction $V$ if and only if ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g$ is. Hence, by the chain rule, e.g. [16, Section 1.4.1] or [5, Theorem 12.2.2], it suffices to show that ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable on all of $(\mathcal{W},\lVert\cdot\rVert_{L_{2}(\mathcal{H})})$ and that $g$ is Gateaux differentiable at $U$ in the direction $V$ in order to show the Gateaux differentiability of $J_{f}=F_{f}\circ\Lambda\circ g$ at $U$ in the direction $V$.
@@
Let $f\in\mathscr{F}$ and $\mathcal{V}\subset\mathcal{H}$ be finite-dimensional. Then $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, i.e. $J_{f}(U_{n})\rightarrow\infty$ whenever $\lVert U_{n}\rVert\rightarrow\infty$. In particular, $J_{f}$ has a global minimum on $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, which can be found among the stationary points of $J_{f}$.
@@
In the proof of Theorem 4.21, coercivity on finite-dimensional subspaces of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the form $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ for finite-dimensional $\mathcal{V}\subset\mathcal{H}$ is enough to show the existence of a global minimizer of $J_{f}$, because all the stationary points lie in one such finite-dimensional subspace.
@@ 4.4 Optimal low-rank posterior covariance approximations
$\min\{D_{\textup{Am},\alpha}(\mathcal{N}(m,\mathcal{C})\|\mu_{\textup{pos}}):\ \mathcal{C}\in\mathscr{C}_{r}\}$ $=D_{\textup{Am},\alpha}(\mathcal{N}(m,\mathcal{C}^{\textup{opt}}_{r}\|\mu_{\textup{pos}}))$
@@ A.1 Hilbert spaces and bounded operators
The proof is a slight modification of the argument of [12, Lemma A.2]. By separability of $\mathcal{H}$ there exists a countable and dense sequence $(h_{i})_{i}$ of $\mathcal{H}$. By density of $\mathcal{D}$ we can construct a countable sequence $(d_{i}^{\prime})_{i}\subset\mathcal{D}$ dense in $\mathcal{H}$ by taking an element of $\mathcal{D}$ from the ball $B(h_{i},1/j)$, for all $i$ and $j\in\mathbb{N}$. Now, we apply Gram–Schmidt to the countable sequence $(e_{1},\ldots,e_{m},d_{1}^{\prime},d_{2}^{\prime},\ldots)\subset\mathcal{D}$ to obtain a countable orthonormal sequence $(d_{i})_{i}\subset\mathcal{D}$. Since $(e_{i})_{i=1}^{m}$ is already orthonormal, $d_{i}=e_{i}$ for $i\leq m$. Furthermore, $d_{i}^{\prime}\in\operatorname{span}{\left(d_{j},j\leq m+i\right)}$. It follows that $(d_{i}^{\prime})_{i}\subset\operatorname{span}{\left((d_{i})_{i}\right)}$ and since $(d_{i}^{\prime})_{i}$ is dense, so is $\operatorname{span}{\left((d_{i})_{i}\right)}$. ∎
@@ B.1 Proofs of Section 3
By multiplying (23) by $A^{-1}$, $A^{-1}(A^{-1})^{*}A^{-1}$ and $(A^{-1})^{*}A^{-1}$ and rearranging, we obtain respectively,
@@
Therefore, $(\tfrac{-\lambda_{i}}{1+\lambda_{i}})_{i\in\mathbb{N}}$ contains exactly $\operatorname{rank}\left({H}\right)\leq n$ many nonzero entries. It follows directly from (10a), (10b) and the fact that $\lambda_{i}\not=0$ for $i\leq\operatorname{rank}\left({H}\right)$, that $w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({H}\right)$. By Lemma A.1, we can extend $(w_{i})_{i=1}^{\operatorname{rank}\left({H}\right)}$ to an ONB $(w_{i}^{\prime})_{i}$ of $\mathcal{H}$ with $(w_{i}^{\prime})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $w_{i}^{\prime}=w_{i}$ for $i\leq\operatorname{rank}\left({H}\right)$. We now replace $w_{i}$ by $w_{i}^{\prime}$ and $v_{i}$ by $\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}^{\prime}$ for $i>\operatorname{rank}\left({H}\right)$. The equations (10a) and (10b) and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for all $i$ remain to hold, and we now have $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$.
@@
Multiplying both sides of the equation from the left by $\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}$ yields (10c).

### v4

@@ 1.4 Outline
We introduce key notation in Section 1.5 below. In Section 2, we begin by recalling the infinite-dimensional formulation of the linear Gaussian inverse problem, and formulate the posterior covariance and posterior precision approximation classes that define our measure approximation problems. In Section 3 we recall the Feldman–Hajek theorem, which characterises when two Gaussian measures are equivalent, and recall expressions for the Kullback–Leibler divergence and Rényi divergence of two equivalent Gaussians. We state the first main result of this paper, Proposition 3.7, which identifies the generalised eigenpairs of the three operator pencils mentioned in the introduction and identifies the finite-dimensional subspace on which the posterior differs from the prior. In Section 4, we consider measure approximation problems where the posterior covariance is approximated and identify solutions in Theorem 4.21 and Corollary 4.23. Auxiliary results are presented in Appendix A, and proofs of the results in this work can be found in Appendix B.
@@ 1.5 Notation
For $I$ a non-empty interval in $\mathbb{R}$, $\ell^{2}(I)$ denotes the space of square-summable sequences, i.e. $\ell^{2}(I)=\{(x_{i})_{i\in\mathbb{N}}\subset I\ :\ \sum_{i\in\mathbb{N}}\lvert x_{i}\rvert^{2}<\infty\}$. If $I\subset\mathbb{R}$ is open, then $C^{1}(I)$ denotes the set of continuously differentiable functions on $I$.
@@ 2 Low-rank posterior covariance approximations
see e.g. [32, Example 6.23]. Equation (3c) should be understood to imply the following two facts: $\operatorname{ran}{\mathcal{C}}_{\textup{pos}}\coloneqq\operatorname{dom}{\mathcal{C}_{\textup{pr}}^{-1}+H}=\operatorname{ran}{\mathcal{C}}_{\textup{pr}}$, and $\mathcal{C}_{\textup{pr}}^{-1}+H:\operatorname{ran}{\mathcal{C}_{\textup{pr}}}\rightarrow\mathcal{H}$ is the inverse of the operator $\mathcal{C}_{\textup{pos}}$ given in (3b). While all nondegenerate Gaussians are equivalent in a finite-dimensional setting, this is no longer true in an infinite-dimensional setting, where in fact it holds that nondegenerate Gaussians that are not equivalent must be mutually singular. By [32, Theorem 6.31], $\mu_{\textup{pos}}$ and $\mu_{\textup{pr}}$ are in fact equivalent. In particular, $\mu_{\textup{pos}}$ is a nondegenerate measure and the above properties of $\mathcal{C}_{\textup{pr}}$ also hold for $\mathcal{C}_{\textup{pos}}$. We shall construct approximations to $\mu_{\textup{pos}}$ that are equivalent to $\mu_{\textup{pos}}$.
@@ 4.2 Equivalence to target measures of low-rank Gaussian approximations
A natural generalisation to infinite dimensions of the low-rank approximation classes for covariance and precision considered in the finite-dimensional setting of [31, eqs. (2.4) and (4.1)], is to take $\widetilde{\mathscr{C}}_{r}\coloneqq\{\mathcal{C}_{\textup{pr}}-KK^{*}>0:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$ and $\widetilde{\mathscr{P}}_{r}=\{\mathcal{C}_{\textup{pr}}^{-1}+UU^{*}:\ U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$. Let $\mathscr{C}_{r,\mathcal{E}}\coloneqq\{\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathcal{E}:\ K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\}$ as in the start of Section 4.2. By the preceding discussion, we have the proper inclusions $\widetilde{\mathscr{P}}_{r}^{-1}\subset\mathscr{C}_{r,\mathcal{E}}$, and by the equivalence (iii)(b)$\Leftrightarrow$(iii)(a) of Lemma 4.5(iii) also the proper inclusion $\mathscr{C}_{r,\mathcal{E}}\subset\widetilde{\mathscr{C}}_{r}$ holds. That is, in general $\widetilde{\mathscr{P}}_{r}^{-1}$ contains strictly fewer covariances than those that maintain equivalence between the resulting approximate and exact posterior, while $\widetilde{\mathscr{C}}_{r}$ contains strictly more. For the loss classes considered in this work which require densities to exist, the set $\widetilde{\mathscr{C}}_{r}$ is thus not suitable. Note that by the definition of $\mathscr{P}_{r}$ in (5), $\widetilde{\mathscr{P}}_{r}=\mathscr{P}_{r}$. The fact that $\widetilde{\mathscr{P}}_{r}^{-1}\subset\mathscr{C}_{r,\mathcal{E}}$ motivates the use of approximation class $\widetilde{\mathscr{P}}_{r}$ in this work. Because the form of the precision updates in $\widetilde{\mathscr{P}}_{r}$ parallels the form in [31, eq. (4.1)] that lies at the core of the approximation procedure of [31], our work can be considered to naturally generalise [31].
@@ 4.3 Differentiability and minimisers of covariance loss function
In order to solve Problem 4.4, we formulate it as a minimisation problem over the set of $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. By Corollary 4.8(ii), solving Problem 4.3 is equivalent to solving Problem 4.4. We want to find the minimiser of the function
@@
We finally define the function $g:\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\rightarrow L_{2}(\mathcal{H})_{\mathbb{R}}$ by
@@
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. If $\mathcal{W}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$ is a subspace of finite dimension that contains $g(U+tV)$ for all $t\in\mathbb{R}$, then the restriction ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}:\mathcal{W}\rightarrow\mathbb{R}$ of $F_{f}\circ\Lambda$ to $\mathcal{W}$ satisfies ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)=F_{f}\circ\Lambda\circ g(U+tV)$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g$ is Gateaux differentiable at $U$ in the direction $V$ if and only if ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g$ is. Hence, by the chain rule, e.g. [16, Section 1.4.1] or [5, Theorem 12.2.2], it suffices to show that ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable on all of $(\mathcal{W},\lVert\cdot\rVert_{L_{2}(\mathcal{H})})$ and that $g$ is Gateaux differentiable at $U$ in the direction $V$ in order to show the Gateaux differentiability of $J_{f}=F_{f}\circ\Lambda\circ g$ at $U$ in the direction $V$. This observation is useful, since such a finite-dimensional subspace $\mathcal{W}$ exists, e.g. $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H}):\ \operatorname{ran}{X}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}+\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}}+\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}VV^{*}\mathcal{C}_{\textup{pos}}^{1/2}}+\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}$}. This subspace is finite-dimensional because $U$, $V$, and $H$ are finite-rank.
@@
Let $f\in\mathscr{F}$ and $\mathcal{V}\subset\mathcal{H}$ be finite-dimensional. Then $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, i.e. $J_{f}(U_{n})\rightarrow\infty$ whenever $\lVert U_{n}\rVert\rightarrow\infty$. In particular, $J_{f}$ has a global minimum on $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, which can be found among the stationary points of the restriction of $J_{f}$ to $\mathcal{B}(\mathbb{R}^{r},\mathcal{V}$).
@@
In the proof of Theorem 4.21, coercivity on finite-dimensional subspaces of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the form $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ for finite-dimensional $\mathcal{V}\subset\mathcal{H}$ is enough to show the existence of a global minimiser of $J_{f}$, because all the stationary points lie in one such finite-dimensional subspace.
@@ 4.4 Optimal low-rank posterior covariance approximations
$\min\{D_{\textup{Am},\alpha}(\mathcal{N}(m,\mathcal{C})\|\mu_{\textup{pos}}):\ \mathcal{C}\in\mathscr{C}_{r}\}$ $=D_{\textup{Am},\alpha}(\mathcal{N}(m,\mathcal{C}^{\textup{opt}}_{r})\|\mu_{\textup{pos}})$
@@ A.1 Hilbert spaces and bounded operators
The proof is a slight modification of the argument of [12, Lemma A.2]. By separability of $\mathcal{H}$ there exists a countable and dense sequence $(h_{i})_{i}$ of $\mathcal{H}$. By density of $\mathcal{D}$ we can construct a countable sequence $(d_{i}^{\prime})_{i}\subset\mathcal{D}$ that is dense in $\mathcal{H}$ by taking an element of $\mathcal{D}$ from the ball $B(h_{i},1/j)$, for all $i$ and $j\in\mathbb{N}$. Now, we apply Gram–Schmidt to the countable sequence $(e_{1},\ldots,e_{m},d_{1}^{\prime},d_{2}^{\prime},\ldots)\subset\mathcal{D}$ to obtain a countable orthonormal sequence $(d_{i})_{i}\subset\mathcal{D}$. Since $(e_{i})_{i=1}^{m}$ is already orthonormal, $d_{i}=e_{i}$ for $i\leq m$. Furthermore, $d_{i}^{\prime}\in\operatorname{span}{\left(d_{j},j\leq m+i\right)}$. It follows that $(d_{i}^{\prime})_{i}\subset\operatorname{span}{\left((d_{i})_{i}\right)}$ and since $(d_{i}^{\prime})_{i}$ is dense, so is $\operatorname{span}{\left((d_{i})_{i}\right)}$. ∎
@@ B.1 Proofs of Section 3
By applying $A^{-1}$, $A^{-1}(A^{-1})^{*}A^{-1}$ and $(A^{-1})^{*}A^{-1}$ to (23) and rearranging, we obtain respectively,
@@
Therefore, $(\tfrac{-\lambda_{i}}{1+\lambda_{i}})_{i\in\mathbb{N}}$ contains exactly $\operatorname{rank}\left({H}\right)\leq n$ many nonzero entries. It follows directly from (10a), (10b) and the fact that $\lambda_{i}\not=0$ for $i\leq\operatorname{rank}\left({H}\right)$, that $w_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for $i\leq\operatorname{rank}\left({H}\right)$. By Lemma A.1, we can extend $(w_{i})_{i=1}^{\operatorname{rank}\left({H}\right)}$ to an ONB $(w_{i}^{\prime})_{i}$ of $\mathcal{H}$ with $(w_{i}^{\prime})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $w_{i}^{\prime}=w_{i}$ for $i\leq\operatorname{rank}\left({H}\right)$. We now replace $w_{i}$ by $w_{i}^{\prime}$ and $v_{i}$ by $\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}^{\prime}$ for $i>\operatorname{rank}\left({H}\right)$. After this replacement, the equations (10a) and (10b) and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for all $i$ remain valid, and we now have $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ for all $i$.
@@
Applying $\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}$ to both sides of the equation yields (10c).